OpenStudy (anonymous):

Josh hit a baseball, and the flight of the ball was modeled by the following equation: h = -16t2 + 65t + 3. After how many seconds will the ball fall to the ground if it is not caught? Round your answer to the nearest tenth. A. 4.0 seconds B. 3.9 seconds C. 4.1 seconds D. 3.8 seconds

OpenStudy (anonymous):

-16t2 + 65t + 3 = 0 solve and positive root will give you the answer

OpenStudy (anonymous):

C

OpenStudy (chaise):

h = -16t^2 + 65t + 3 When the height is zero the ball will be on the ground. Once the equation is equal to zero we can solve this equation. When we solve it we will get a positive and negative answer, the negative answer will be assumed negligent as in this circumstance it is not possible for time to be negative. 0 = -16t^2 + 65t +3 The quadratic formula is: t = (-b + sqrt(b^2-4ac)/2a For some reason my equation editor isn't working. Sorry. In this equation: a = -16 b = 65 c = 3 I'm not going to show all steps in the quadratic formula, however, if you substitute these values. You should get: t = (1/32)(65+sqrt(4417)) t = 4.108 The ball will take approx. 4.1 seconds to reach the ground if it is not caught.

OpenStudy (anonymous):

good work thanks for showing workings

OpenStudy (chaise):

My pleasure (: If you need me to show working out for the quadratic formula I would be happy to help.

OpenStudy (anonymous):

i will on my next one..